OpenStudy (anonymous):

y=2^x+1 find the y-intercept as an ordered pair and find the range

OpenStudy (anonymous):



OpenStudy (lgbasallote):

\[y = 2^{x+1}\] to find the y-intercept let x = 0 then solve for y

OpenStudy (lgbasallote):

wait...my mistake

OpenStudy (lgbasallote):

\[\huge y = 2^x + 1\] to find y-intercept let x = 0 then solve fory

OpenStudy (lgbasallote):

for y*

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the small x^ is confusing to me

OpenStudy (lgbasallote):

if you let x = 0 it will become \[\huge y = 2^0 + 1\] right?

OpenStudy (anonymous):

ok

OpenStudy (lgbasallote):

...ok?

OpenStudy (anonymous):

so when you have a 0 ^ how does that effect the 2?

OpenStudy (anonymous):

does it leave it as a 2 or make it a 0

OpenStudy (anonymous):

it becomes a 1

OpenStudy (anonymous):

anything to the power of 0 becomes a 1

OpenStudy (anonymous):

so y=2

OpenStudy (anonymous):

yeah y=2 when x=0 put those values into order pair form (x,y)

OpenStudy (anonymous):

(0,2)....how do I find the y intercept

OpenStudy (lgbasallote):

that is the y-intercept...

OpenStudy (anonymous):

sorry I meant the range

OpenStudy (lgbasallote):

to find the range, you start by solving for x..do you know how to? or do you need more help?

OpenStudy (anonymous):

so if I make x=1 then y=3 (1,3)

OpenStudy (lgbasallote):

why would you make x = 1?

OpenStudy (anonymous):

what do you mean solving for x?

OpenStudy (lgbasallote):

you make the equation look like x = need help?

OpenStudy (anonymous):

I don't know how to do that when it is a "^"